OpenStudy (anonymous):

sqrt180 - sqrt125 + sqrt5

Nnesha (nnesha):

\[\huge\rm \sqrt{180}- \sqrt{125}+\sqrt{5}\] factor 125 and 180

OpenStudy (anonymous):

16750 I believe

OpenStudy (anonymous):

whoops did it wrong

OpenStudy (anonymous):

16800 is the answer

Nnesha (nnesha):

what are factors of 125 ?

OpenStudy (anonymous):

1,5,25,125

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha what are factors of 125 ? \(\color{blue}{\text{End of Quote}}\) this question is for @JodiChicken

Nnesha (nnesha):

:-)

Nnesha (nnesha):

okay so factor of 125 are 1,5,25,125 we have to pick one number that should be perfect square

OpenStudy (anonymous):

Ik sorry

Nnesha (nnesha):

it's okay :-)

OpenStudy (anonymous):

125 is 5,5,5,. 180 is 3,3,2,2,5. Then you have for 180,an index of 6, radicand 5. 125 woollen index 5, radicand 5.

OpenStudy (anonymous):

@Nnesha he/she answered

Nnesha (nnesha):

factor of 180 aare 1,2,3,4,5,6,9,10,12,15,18,20,30,36,45,60,90,180 we need greatest factor which should BE perfect square root (meaning not decimal)

Nnesha (nnesha):

yes right 125 = 5 times 5 times 5 so \[\huge\rm \sqrt{180}- \sqrt{5 \times 5^2}+\sqrt{5}\] square of 5 can cancel out with square root so we can write it outside the root so \[\huge\rm \sqrt{36 \times 5}- 5\sqrt{5 }+\sqrt{5}\] greatest factor of 180 is 36 36 times 5=180 take square root of 36 let me know what you get